# A complete characterization of testable hypotheses

### v1

@@ Abstract
We revisit a fundamental question in hypothesis testing: given two sets of probability measures $\mathcal{P}$ and $\mathcal{Q}$, when does a nontrivial (i.e. strictly unbiased) test for $\mathcal{P}$ against $\mathcal{Q}$ exist? Le Cam showed that, when $\mathcal{P}$ and $\mathcal{Q}$ have a common dominating measure, a test that has power exceeding its level by more than $\varepsilon$ exists if and only if the convex hulls of $\mathcal{P}$ and $\mathcal{Q}$ are separated in total variation distance by more than $\varepsilon$. The requirement of a dominating measure is frequently violated in nonparametric statistics. In a passing remark, Le Cam described an approach to address more general scenarios, but he stopped short of stating a formal theorem. This work completes Le Cam’s program, by presenting a matching necessary and sufficient condition for testability: for the aforementioned theorem to hold without assumptions, one must take the closures of the convex hulls of $\mathcal{P}$ and $\mathcal{Q}$ in the space of bounded finitely additive measures. We provide simple elucidating examples, and elaborate on various subtle measure theoretic and topological points regarding compactness and achievability.
@@ 1 Introduction
The minimax risk is closely related to the total variation (TV) distance between $\operatorname{co}({\mathcal{P}})$ and $\operatorname{co}({\mathcal{Q}})$. Indeed, for any test $\phi$ and arbitrary probability measures $\mu,\nu$, it straightforwardly holds that
@@
In words, if $\mathcal{P}$} and $\mathcal{Q}$} have a common dominating measure, then a nontrivial test exists if and only if the convex hulls of $\mathcal{P}$} and $\mathcal{Q}$} are separated in TV distance. Actually, Kraft (1955) presented only (1.3) and proved it using the Hahn–Banach separating hyperplane theorem. A statement like (1.4) appears on page 476 of Le Cam (2012) and was proved succinctly using a minimax theorem, but the left hand side there is subtly (yet critically!) different; we return to Le Cam’s claim in Theorem 1.11 and later again in Section 4. Also using a Hahn–Banach argument, Proposition 1.6 shows that Theorem 1.1 is a corollary of our own more general Theorem 1.5, itself proved using a minimax theorem.
@@
The infimum TV distance is not achieved in general in Theorem 1.1. To see this, let ${\mathcal{Q}}=\{\delta_{0}\}$ and ${\mathcal{P}}=\{(\frac{1}{2}-\frac{1}{n})\delta_{0}+(\frac{1}{2}+\frac{1}{n})\delta_{n}\colon n\in{\mathbb{N}}\}$. Then $\operatorname{d}_{\textnormal{TV}}(\operatorname{co}({\mathcal{P}}),\delta_{0})=1/2$, but this is not achieved by any $\mu\in\operatorname{\overline{co}}({\mathcal{P}})$ because there is no such $\mu$ with mass exactly 1/2 at 0. Said differently, the TV projection of a measure onto $\operatorname{\overline{co}}({\mathcal{P}})$ may not exist. (In fact, there will always exist a projection onto $\operatorname{\overline{co}*}({\mathcal{P}})$, an object we introduce later.)
@@
Let ${\mathcal{P}}=\{\delta_{x}:x\in[0,1]\}$, where $\delta_{x}$ is the Dirac delta mass at $x$, and let ${\mathcal{Q}}=\{\mathrm{Uni}[0,1]\}$, where $\mathrm{Uni}[0,1]$ denotes the uniform distribution on $[0,1]$. The TV distance between $\mathrm{Uni}[0,1]$ and $\operatorname{\overline{co}}({\mathcal{P}})$ equals 1, the largest possible value. However, as should be intuitively clear, and also follows as a consequence of our main theorem below, no nontrivial test exists for this problem, and the minimax risk is $R({\mathcal{P}},{\mathcal{Q}})=1$. [^4] Note that, as stated, in this example any test $\phi$ maps a single observation to $[0,1]$. Letting ${\mathcal{P}}=\{\delta_{x}\times\delta_{x}:x\in[0,1]\}$ and ${\mathcal{Q}}=\{\mathrm{Uni}[0,1]\times\mathrm{Uni}[0,1]\}$, one can design a perfect test that has zero type-I error and power one (the test rejects if the two observations are different). While our setup allows the underlying $\Omega$ and distributions to be arbitrary, and covers the case of $\mathcal{P}$} and $\mathcal{Q}$} consisting of product distributions as a special case, the purpose of the example is to show that Theorem 1.1 can fail in nondominated settings. (See also Proposition A.1.)
@@
In fact,
$R({\mathcal{P}},{\mathcal{Q}})=1-\operatorname{d}_{\textnormal{TV}}(\operatorname{\overline{co}*}({\mathcal{P}}),\operatorname{\overline{co}*}({\mathcal{Q}})).$ (1.5)
@@
We suspect that outside the dominated setting of Theorem 1.1, the minimax risk may not in general be achieved by any test $\phi$, but we do not currently have a counterexample that confirms this.
@@
The above result holds despite the fact that $\operatorname{\overline{co}}({\mathcal{P}})\neq\operatorname{\overline{co}*}({\mathcal{P}})$ in general. To see that the two closures can be different, we may return to the example where ${\mathcal{P}}=\{\delta_{n}\colon n\in{\mathbb{N}}\}$ and note that $\operatorname{\overline{co}*}({\mathcal{P}})$ includes a purely finitely additive measure that is obtained as a limit point of the sequence of point masses $\delta_{n}$. We denote this limit point by $\delta_{\infty}$ because it is “supported at infinity”, in that it assigns zero mass to any finite subset of $\mathbb{N}$}. (We may also note this from Example 1.2, where $\operatorname{\overline{co}*}({\mathcal{P}})$ contains an equal mixture of $\delta_{0}$ and $\delta_{\infty}$.)
@@
In particular, (1.3) holds and the test $\phi$ can be taken to be continuous therein. The infimum in the total-variation distance $d_{\textnormal{TV}}$ is achieved by some $\mu\in{\mathcal{P}},\nu\in{\mathcal{Q}}$.
@@
Example 1.8.
Let us first note that in general, ${\mathcal{P}}\subsetneq\operatorname{\overline{co}*}({\mathcal{P}})$. Consider $\Omega=[-1,1]$ and $\mathcal{P}$} to be the set of all symmetric laws around 0, which is convex and weakly compact. Now, consider the sequence $\mu_{n}=\mathrm{Uni}[-1/n,1/n]$, which converges weakly to $\delta_{0}$, but does not converge to $\delta_{0}$ in the weak-* topology. (Indeed, the test function $\phi(x)=0$ for $x\neq 0$ and $\phi(x)=1$ for $x=0$ satisfies ${\mathbb{E}}_{\mu_{n}}[\phi]=0$, but ${\mathbb{E}}_{\delta_{0}}[\phi]=1$, and such discontinuous test functions are permissible in the weak-* topology.) Instead, the sequence $(\mu_{n})$ has a subnet that converges to an element in ba, showing that ${\mathcal{P}}\subsetneq\operatorname{\overline{co}*}({\mathcal{P}})$ as claimed. Despite this, the TV distance in the above theorem is unaffected by taking the closure in ba because in a metric space with the Borel sigma algebra, $\operatorname{d}_{\textnormal{TV}}(\mu,\nu)=\sup_{\phi\in\Phi\cap C_{b}}({\mathbb{E}}_{\mu}[\phi]-{\mathbb{E}}_{\nu}[\phi])$. That is, we only need to consider continuous $[0,1]$-bounded functions to compute the TV distance; see Lemma 2.3. In fact, this lemma is the only place where the stated assumption on $(\Omega,{\mathcal{F}})$ is used.
@@
Let $\mathcal{P}$} be all probability measures on $[0,1]$ whose mean is at most $m_{1}$, and let $\mathcal{Q}$} be those with mean at least $m_{2}>m_{1}$. These sets do not have a dominating reference measure, but are weakly compact and convex, and so Theorem 1.7 (see below) applies, but we can say more. Consider the test $\phi^{*}(x)=x$ and let $\mu^{*},\nu^{*}$ be Bernoullis with means $m_{1},m_{2}$. Then $R(\phi^{*})$ is easily seen to equal $m_{1}+(1-m_{2})$, which equals $1-\operatorname{d}_{\textnormal{TV}}(\mu^{*},\nu^{*})$. Thus, Corollary 1.9 implies that $\phi^{*}$ is minimax optimal and the two Bernoullis are the TV-closest pair. Thus, here, the minimax optimal test exists, despite Theorem 1.7 not guaranteeing its existence.
The last result that we wish to highlight is that of Huber and Strassen (1973), who prove that if the Choquet capacities induced by ${\mathcal{P}},{\mathcal{Q}}$ are two-alternating, then a least favorable distribution pair exists, such that a likelihood ratio test between them is minimax optimal. This result applies to both parametric settings which may not be compact (e.g., testing unit variance Gaussians with nonpositive mean against those with mean larger than one), as well as nonparametric settings without a reference measure (e.g., when the null and alternative are each described by a total variation ball around a simple hypothesis). Nevertheless, it does not yield a necessary and sufficient condition for testability as Theorem 1.5 does, just a (very useful) sufficient one.
@@
Write $\mathcal{M}$} for the set of finite signed countably additive measures on $\Omega$, which is a Banach space with TV-norm, and let $\mathcal{M}$$^{\prime}$ denote its dual. A generalized test is an element $\phi\in{\mathcal{M}}^{\prime}$ such that $\phi(\mu)\in[0,1]$ for all $\mu\in{\mathcal{M}}_{1}$. (The reader is warned that Le Cam just called these tests, which may lead to confusion unless one very carefully follows his definitions.)
@@
As Le Cam had also noted, these generalized tests may not correspond to bounded measurable functions, and thus need not be possible to evaluate using observed samples. The tests $\phi$ that we consider also live in $\mathcal{M}^{\prime}$, but are in fact given by bounded measurable functions (by definition). Thus the above result, while correct, has limited practical applicability, unlike our main Theorem 1.5. We discuss this point in detail in Section 4.
@@
The rest of this paper is organized as follows. Section 2 presents the mathematical background to understand the weak-* closure in ba, and the proofs of all the results stated above. Section 3 discusses the interesting case of a singleton ${\mathcal{Q}}=\{{\mathsf{Q}}\}$ and its relation to recent work. Section 4 discusses Le Cam’s generalized tests in more detail. Finally, Section 5 proves Theorem 3.4, which is a result relating the closed convex hull in ba to the bipolar of $\mathcal{P}$}, another fundamental object from recent work.
@@ 2 Finitely additive measures and proofs of the main results
A bounded finitely additive measure is a real-valued set function $\mu$ on $\mathcal{F}$} that is (i) additive, $\mu(A\cup B)=\mu(A)+\mu(B)$ for all disjoint sets $A,B\in{\mathcal{F}}$, and (ii) bounded in total variation norm, $\|\mu\|_{\textnormal{TV}}=\sup_{A\in{\mathcal{F}}}|\mu(A)|<\infty$. The space of all such $\mu$ is denoted by ba, the subset taking nonnegative values by $\textnormal{{ba}}_{+}$, and the further subset with unit mass by $\textnormal{{ba}}_{1}$. Thus $\textnormal{{ba}}_{1}$ is the space of finitely additive probability measures, also known as probability charges.
@@
The integral $\int_{\Omega}fd\mu$, or ${\mathbb{E}}_{\mu}[f]$ in probabilistic notation, is defined for any $f\in{\mathcal{L}}$ and any $\mu\in\textnormal{{ba}}$, and is linear in $f$ and $\mu$. Limit theorems such as the dominated and monotone convergence theorems do not apply, but one does have the bound $|{\mathbb{E}}_{\mu}[f]|\leq\|f\|_{\infty}\|\mu\|_{\textnormal{TV}}$ where $\|f\|_{\infty}=\sup_{\omega\in\Omega}|f(\omega)|$ is the supremum norm. This leads to the identification of ba as the dual of the Banach space $({\mathcal{L}},\|\cdot\|_{\infty})$. Like the dual of any Banach space, ba can be equipped with the weak-* topology $\sigma(\textnormal{{ba}},{\mathcal{L}})$, defined as the weakest topology such that the linear functionals $\mu\mapsto{\mathbb{E}}_{\mu}[f]$ are continuous for all $f\in{\mathcal{L}}$. The Banach–Alaoglu theorem implies that the set $\textnormal{{ba}}_{1}$ is weak-* compact, and then so is any weak-* closed subset of $\textnormal{{ba}}_{1}$. In particular, this is the case for $\operatorname{\overline{co}*}({\mathcal{P}})$ and $\operatorname{\overline{co}*}({\mathcal{Q}})$. It is this compactness property under the weak-* topology that explains the appearance of the space ba in Theorem 1.5.
@@
The total variation norm on ba has the representation $\|\mu\|_{\textnormal{TV}}=\sup_{f\in B_{1}}{\mathbb{E}}_{\mu}[f]$, where $B_{1}$ is the unit ball in $\mathcal{L}$}. It is customary to define the total variation distance $\operatorname{d}_{\textnormal{TV}}(\mu,\nu)$ between two probability charges $\mu,\nu$ as half the total variation norm of $\mu-\nu$. Because $\mu,\nu$ have unit mass, and because $B_{1}=2\Phi-1$, meaning that $f\in B_{1}$ if and only if $f=2\phi-1$ for some test $\phi$, one obtains the representation
@@
A similar, but different, calculation can be used to obtain Theorem 1.7. This calculation will make use of the assumption that $\mathcal{P}$} and $\mathcal{Q}$} are convex and weakly compact in the usual sense, i.e., in the topology $\sigma({\mathcal{M}}_{1},C_{b})$. By considering continuous tests $\phi\in\Phi_{c}=\Phi\cap C_{b}$, the map $\mu\mapsto{\mathbb{E}}_{\mu}[\phi]$ remains continuous, and the minimax theorem still applies. Restricting to continuous tests raises the question of whether the formula (2.1) for the total variation distance remains valid with $\Phi_{c}$ in place of $\Phi$. The following lemma confirms that in the metric space setting of Theorem 1.7 this is indeed so.
@@
The Hahn decomposition of the signed measure $\eta=\mu-\nu$ yields a measurable set $A$ such that the positive part $\eta^{+}$ is concentrated on $A$ and the negative part $\eta^{-}$ is concentrated on $A^{c}$. The total variation distance is then $\operatorname{d}_{\textnormal{TV}}(\mu,\nu)=\mu(A)-\nu(A)=\eta^{+}(A)$, the total mass of $\eta^{+}$. Fix any closed sets $F_{1}\subset A$ and $F_{0}\subset A^{c}$. Urysohn’s lemma, which applies in any metric space, yields a continuous function $\phi$ with $\phi=1$ on $F_{1}$, $\phi=0$ on $F_{0}$, and $0\leq\phi\leq 1$. Thus
@@
$\sup_{\phi\in\Phi\cap C_{b}}({\mathbb{E}}_{\mu}[\phi]-{\mathbb{E}}_{\nu}[\phi])\geq\eta^{+}(F_{1})+\eta^{-}(F_{0})-\eta^{-}(A^{c})$
for all closed sets $F_{1}\subset A$ and $F_{0}\subset A^{c}$. Taking the supremum of the right-hand side over all such $F_{1}$ and $F_{0}$ yields $\eta^{+}(A)$, since positive finite measures on a metric space with the Borel $\sigma$-algebra are regular; see Billingsley (1999, Theorem 1.1). We have shown that $\sup_{\phi\in\Phi\cap C_{b}}({\mathbb{E}}_{\mu}[\phi]-{\mathbb{E}}_{\nu}[\phi])\geq\operatorname{d}_{\textnormal{TV}}(\mu,\nu)$, and the reverse inequality follows from the fact that ${\mathbb{E}}_{\mu}[\phi]-{\mathbb{E}}_{\nu}[\phi]\leq\operatorname{d}_{\textnormal{TV}}(\mu,\nu)$ for all $\phi\in\Phi$. ∎
We now proceed with the proof of Theorem 1.7. Although the main calculation is very similar to the one in the proof of Theorem 1.5, we emphasize that it is not subsumed by it. Furthermore, Theorem 1.5 is used in the proof, so the two results are not independent. Recall that $\Phi_{c}=\Phi\cap C_{b}$ is the set of continuous tests.

### v2

@@ Abstract
We revisit a fundamental question in hypothesis testing: given two sets of probability measures $\mathcal{P}$ and $\mathcal{Q}$, when does a nontrivial (i.e. strictly unbiased) test for $\mathcal{P}$ against $\mathcal{Q}$ exist? Le Cam showed that, when $\mathcal{P}$ and $\mathcal{Q}$ have a common dominating measure, a test that has power exceeding its level by more than $\varepsilon$ exists if and only if the convex hulls of $\mathcal{P}$ and $\mathcal{Q}$ are separated in total-variation distance by more than $\varepsilon$. The requirement of a dominating measure is frequently violated in nonparametric statistics. In a passing remark, Le Cam described an approach to address more general scenarios, but he stopped short of stating a formal theorem. This work completes Le Cam’s program, by presenting a matching necessary and sufficient condition for testability: for the aforementioned theorem to hold without assumptions, one must take the closures of the convex hulls of $\mathcal{P}$ and $\mathcal{Q}$ in the space of bounded finitely additive measures. We provide simple elucidating examples, and elaborate on various subtle measure theoretic and topological points regarding compactness and achievability.
@@ 1 Introduction
The minimax risk is closely related to the total-variation (TV) distance between $\operatorname{co}({\mathcal{P}})$ and $\operatorname{co}({\mathcal{Q}})$. Indeed, for any test $\phi$ and arbitrary probability measures $\mu,\nu$, it straightforwardly holds that
@@
In words, if $\mathcal{P}$} and $\mathcal{Q}$} have a common dominating measure, then a nontrivial test exists if and only if the convex hulls of $\mathcal{P}$} and $\mathcal{Q}$} are separated in TV distance. Actually, Kraft (1955) presented only (1.3) and proved it using the Hahn–Banach separating hyperplane theorem. A statement like (1.4) appears on page 476 of Le Cam (2012) and was proved succinctly using a minimax theorem, but the left hand side there is subtly (yet critically!) different; we return to Le Cam’s claim in Theorem 1.11 and later again in Appendix B. Also using a Hahn–Banach argument, Proposition 1.6 shows that Theorem 1.1 is a corollary of our own more general Theorem 1.5, itself proved using a minimax theorem.
@@
The infimum TV distance is not achieved in general in Theorem 1.1. To see this, let ${\mathcal{Q}}=\{\delta_{0}\}$ and ${\mathcal{P}}=\{(\frac{1}{2}-\frac{1}{n})\delta_{0}+(\frac{1}{2}+\frac{1}{n})\delta_{n}\colon n\in{\mathbb{N}}\}$. Then $\operatorname{d}_{\textnormal{TV}}(\operatorname{co}({\mathcal{P}}),\delta_{0})=1/2$, but this is not achieved by any $\mu\in\operatorname{\overline{co}}({\mathcal{P}})$ because there is no such $\mu$ with mass exactly 1/2 at 0. Said differently, the TV projection of a measure onto $\operatorname{\overline{co}}({\mathcal{P}})$ may not exist. (The TV projection will always exist onto $\operatorname{\overline{co}*}({\mathcal{P}})$, an object we introduce later.)
@@
Let ${\mathcal{P}}=\{\delta_{x}:x\in[0,1]\}$, where $\delta_{x}$ is the Dirac delta mass at $x$, and let ${\mathcal{Q}}=\{\mathrm{Uni}[0,1]\}$, where $\mathrm{Uni}[0,1]$ denotes the uniform distribution on $[0,1]$. The TV distance between $\mathrm{Uni}[0,1]$ and $\operatorname{\overline{co}}({\mathcal{P}})$ equals 1, the largest possible value. However, as should be intuitively clear, and also follows as a consequence of our main theorem below, no nontrivial test exists for this problem, and the minimax risk is $R({\mathcal{P}},{\mathcal{Q}})=1$. [^4] Note that, as stated, in this example any test $\phi$ maps a single observation to $[0,1]$. Letting ${\mathcal{P}}=\{\delta_{x}\times\delta_{x}:x\in[0,1]\}$ and ${\mathcal{Q}}=\{\mathrm{Uni}[0,1]\times\mathrm{Uni}[0,1]\}$, one can design a perfect test that has zero type-I error and power one (the test rejects if the two observations are different). While our setup allows the underlying $\Omega$ and distributions to be arbitrary, and covers the case of $\mathcal{P}$} and $\mathcal{Q}$} consisting of product distributions as a special case, the purpose of the example is to show that Theorem 1.1 can fail in nondominated settings. (See also Proposition C.1.)
@@
In fact, we have
$R({\mathcal{P}},{\mathcal{Q}})=1-\operatorname{d}_{\textnormal{TV}}(\operatorname{\overline{co}*}({\mathcal{P}}),\operatorname{\overline{co}*}({\mathcal{Q}})),$ (1.5)
@@
We suspect that outside of special cases, such as the dominated setting of Theorem 1.1, the minimax risk may not in general be achieved by any test $\phi$.
The proof of the above result is based on a minimax theorem, as detailed in the following section. We note that the choice of topology is a critical and subtle matter in this regard, because it affects both the compactness of sets and continuity of functions, required for invoking an appropriate minimax theorem. In Appendix A we expand briefly on this, noting that the established literature contains instances where minimax theorems are applied incorrectly.
@@
The above result holds despite the fact that $\operatorname{\overline{co}}({\mathcal{P}})\neq\operatorname{\overline{co}*}({\mathcal{P}})$ in general. To see that the two closures can be different, we an example where ${\mathcal{P}}=\{\delta_{n}\colon n\in{\mathbb{N}}\}$ and note that $\operatorname{\overline{co}*}({\mathcal{P}})$ includes a purely finitely additive measure that is obtained as a limit point of the sequence of point masses $\delta_{n}$. We denote this limit point by $\delta_{\infty}$ because it is “supported at infinity”, in that it assigns zero mass to any finite subset of $\mathbb{N}$}. (We may also note this from Example 1.2, where $\operatorname{\overline{co}*}({\mathcal{P}})$ contains an equal mixture of $\delta_{0}$ and $\delta_{\infty}$.)
@@
In particular, (1.3) holds and the test $\phi$ can be taken to be continuous therein, and the infimum in the total-variation distance $d_{\textnormal{TV}}$ is achieved by some $\mu\in{\mathcal{P}},\nu\in{\mathcal{Q}}$.
@@
Remark 1.8.
In the context of Theorem 1.7, let us first note that in general, ${\mathcal{P}}\subsetneq\operatorname{\overline{co}*}({\mathcal{P}})$. Consider $\Omega=[-1,1]$ and $\mathcal{P}$} to be the set of all symmetric laws around 0, which is convex and weakly compact. Now, consider the sequence $\mu_{n}=\mathrm{Uni}[-1/n,1/n]$, which converges weakly to $\delta_{0}$, but does not converge to $\delta_{0}$ in the weak-* topology. (Indeed, the test function $\phi(x)=0$ for $x\neq 0$ and $\phi(x)=1$ for $x=0$ satisfies ${\mathbb{E}}_{\mu_{n}}[\phi]=0$, but ${\mathbb{E}}_{\delta_{0}}[\phi]=1$, and such discontinuous test functions are permissible in the weak-* topology.) Instead, the sequence $(\mu_{n})$ has a subnet that converges to an element in ba, showing that ${\mathcal{P}}\subsetneq\operatorname{\overline{co}*}({\mathcal{P}})$ as claimed. Despite this, the TV distance in the above theorem is unaffected by taking the closure in ba because in a metric space with the Borel sigma algebra, $\operatorname{d}_{\textnormal{TV}}(\mu,\nu)=\sup_{\phi\in\Phi\cap C_{b}}({\mathbb{E}}_{\mu}[\phi]-{\mathbb{E}}_{\nu}[\phi])$; see Lemma 2.3. That is, we only need to consider continuous $[0,1]$-bounded functions to compute the TV distance. In fact, Lemma 2.3 is the only reason we need a metric, rather than merely a topology, in Theorem 1.7.
@@
Let $\mathcal{P}$} be all probability measures on $[0,1]$ whose mean is at most $m_{1}$, and let $\mathcal{Q}$} be those with mean at least $m_{2}>m_{1}$. These sets do not have a dominating reference measure, but are weakly compact and convex, and so Theorem 1.7 applies, but we can say more. Consider the test $\phi^{*}(x)=x$ and let $\mu^{*},\nu^{*}$ be Bernoullis with means $m_{1},m_{2}$. Then $R(\phi^{*})$ is easily seen to equal $m_{1}+(1-m_{2})$, which equals $1-\operatorname{d}_{\textnormal{TV}}(\mu^{*},\nu^{*})$. Thus, Corollary 1.9 implies that $\phi^{*}$ is minimax optimal and the two Bernoullis are the TV-closest pair. Thus, here, the minimax optimal test exists, despite Theorem 1.7 not guaranteeing its existence.
The last result that we wish to highlight is that of Huber and Strassen (1973), who prove that if the Choquet capacities induced by ${\mathcal{P}},{\mathcal{Q}}$ are two-alternating, then a least favorable distribution pair exists, such that a likelihood ratio test between them is minimax optimal. This result applies to both parametric settings which may not be compact (e.g., testing unit variance Gaussians with nonpositive mean against those with mean larger than one), as well as nonparametric settings without a reference measure (e.g., when the null and alternative are each described by a total-variation ball around a simple hypothesis). Nevertheless, it does not yield a necessary and sufficient condition for testability as Theorem 1.5 does, just a (very useful) sufficient one.
@@
Write $\mathcal{M}$} for the set of finite signed countably additive measures on $\Omega$, which is a Banach space when equipped with the TV-norm, and let $\mathcal{M}$$^{\prime}$ denote its dual. A generalized test is an element $\phi\in{\mathcal{M}}^{\prime}$ such that $\phi(\mu)\in[0,1]$ for all $\mu\in{\mathcal{M}}_{1}$. (The reader is warned that Le Cam just called these tests, which may lead to confusion unless one very carefully follows his definitions.)
@@
As Le Cam had also noted, these generalized tests may not correspond to bounded measurable functions, and thus need not be possible to evaluate using observed samples. The tests $\phi$ that we consider also live in $\mathcal{M}^{\prime}$, but are in fact given by bounded measurable functions (by definition). Thus the above result, while correct, has limited practical applicability, unlike our main Theorem 1.5. We discuss this point in detail in Appendix B.
@@
The rest of this paper is organized as follows. Section 2 presents the mathematical background to understand the weak-* closure in ba, and the proofs of all the results stated above. Section 3 discusses the interesting case of a singleton ${\mathcal{Q}}=\{\nu\}$ and its relation to recent work. Appendix A elaborates on some subtle issues related to the choice of topology. Appendix B discusses Le Cam’s generalized tests in more detail. Appendix C contains some auxiliary technical results. Finally, Appendix D proves Theorem 3.4, which is a result relating the closed convex hull in ba to the bipolar of $\mathcal{P}$}, another fundamental object from recent work.
@@ 2 Finitely additive measures and proofs of the main results
A bounded finitely additive measure is a real-valued set function $\mu$ on $\mathcal{F}$} that is (i) additive, $\mu(A\cup B)=\mu(A)+\mu(B)$ for all disjoint sets $A,B\in{\mathcal{F}}$, and (ii) bounded in total-variation norm, $\|\mu\|_{\textnormal{TV}}=\sup_{A\in{\mathcal{F}}}|\mu(A)|<\infty$. The space of all such $\mu$ is denoted by ba, the subset taking nonnegative values by $\textnormal{{ba}}_{+}$, and the further subset with unit mass by $\textnormal{{ba}}_{1}$. Thus $\textnormal{{ba}}_{1}$ is the space of finitely additive probability measures, also known as probability charges.
@@
The integral $\int_{\Omega}fd\mu$, or ${\mathbb{E}}_{\mu}[f]$ in probabilistic notation, is defined for any $f\in{\mathcal{L}}$ and any $\mu\in\textnormal{{ba}}$, and is linear in $f$ and $\mu$. Limit theorems such as the dominated and monotone convergence theorems do not apply, but one does have the bound $|{\mathbb{E}}_{\mu}[f]|\leq\|f\|_{\infty}\|\mu\|_{\textnormal{TV}}$, where $\|f\|_{\infty}=\sup_{\omega\in\Omega}|f(\omega)|$ is the supremum norm. This leads to the identification of ba as the dual of the Banach space $({\mathcal{L}},\|\cdot\|_{\infty})$. Like the dual of any Banach space, ba can be equipped with the weak-* topology $\sigma(\textnormal{{ba}},{\mathcal{L}})$, defined as the weakest topology such that the linear functionals $\mu\mapsto{\mathbb{E}}_{\mu}[f]$ are continuous for all $f\in{\mathcal{L}}$. The Banach–Alaoglu theorem implies that the set $\textnormal{{ba}}_{1}$ is weak-* compact, and then so is any weak-* closed subset of $\textnormal{{ba}}_{1}$. In particular, this is the case for $\operatorname{\overline{co}*}({\mathcal{P}})$ and $\operatorname{\overline{co}*}({\mathcal{Q}})$. It is this compactness property under the weak-* topology that explains the appearance of the space ba in Theorem 1.5.
@@
The total-variation norm on ba has the representation $\|\mu\|_{\textnormal{TV}}=\sup_{f\in B_{1}}{\mathbb{E}}_{\mu}[f]$, where $B_{1}$ is the unit ball in $\mathcal{L}$}. It is customary to define the total-variation distance $\operatorname{d}_{\textnormal{TV}}(\mu,\nu)$ between two probability charges $\mu,\nu$ as half the total-variation norm of $\mu-\nu$. Because $\mu,\nu$ have unit mass, and because $B_{1}=2\Phi-1$, meaning that $f\in B_{1}$ if and only if $f=2\phi-1$ for some test $\phi$, one obtains the representation
@@
A similar, but different, calculation can be used to obtain Theorem 1.7. This calculation will make use of the assumption that $\mathcal{P}$} and $\mathcal{Q}$} are convex and weakly compact in the usual sense, i.e., in the topology $\sigma({\mathcal{M}}_{1},C_{b})$. By considering continuous tests $\phi\in\Phi_{c}=\Phi\cap C_{b}$, the map $\mu\mapsto{\mathbb{E}}_{\mu}[\phi]$ remains continuous, and the minimax theorem still applies. Restricting to continuous tests raises the question of whether the representation in (2.1) for the total-variation distance remains valid with $\Phi_{c}$ in place of $\Phi$. The following lemma confirms that in the metric space setting of Theorem 1.7 this is indeed so.
@@
The Hahn decomposition of the signed measure $\eta=\mu-\nu$ yields a measurable set $A$ such that the positive part $\eta^{+}$ is concentrated on $A$ and the negative part $\eta^{-}$ is concentrated on $A^{c}$. The total-variation distance is then $\operatorname{d}_{\textnormal{TV}}(\mu,\nu)=\mu(A)-\nu(A)=\eta^{+}(A)$, the total mass of $\eta^{+}$. Fix any closed sets $F_{1}\subset A$ and $F_{0}\subset A^{c}$. Urysohn’s lemma, which applies in any metric space, yields a continuous function $\phi$ with $\phi=1$ on $F_{1}$, $\phi=0$ on $F_{0}$, and $0\leq\phi\leq 1$. Thus
@@
$\sup_{\phi\in\Phi_{c}}({\mathbb{E}}_{\mu}[\phi]-{\mathbb{E}}_{\nu}[\phi])\geq\eta^{+}(F_{1})+\eta^{-}(F_{0})-\eta^{-}(A^{c})$
for all closed sets $F_{1}\subset A$ and $F_{0}\subset A^{c}$. Taking the supremum of the right-hand side over all such $F_{1}$ and $F_{0}$ yields $\eta^{+}(A)$, since positive finite measures on a metric space with the Borel $\sigma$-algebra are regular; see Billingsley (1999, Theorem 1.1). We have shown that $\sup_{\phi\in\Phi_{c}}({\mathbb{E}}_{\mu}[\phi]-{\mathbb{E}}_{\nu}[\phi])\geq\operatorname{d}_{\textnormal{TV}}(\mu,\nu)$, and the reverse inequality follows from the fact that ${\mathbb{E}}_{\mu}[\phi]-{\mathbb{E}}_{\nu}[\phi]\leq\operatorname{d}_{\textnormal{TV}}(\mu,\nu)$ for all $\phi\in\Phi$. ∎
We now proceed with the proof of Theorem 1.7. Although the main calculation is very similar to the one in the proof of Theorem 1.5, we emphasize that it is not subsumed by it. Furthermore, Theorem 1.5 is used in the proof, so the two results are not independent. Recall that $\Phi_{c}=\Phi\cap C_{b}$ denotes the set of continuous tests.